OpenStudy (anonymous):

Fin the area of a rectangle with lenght (5x-15)/(4) cm and area (x^2-5x+6)cm^2

OpenStudy (ybarrap):

(Assuming you meant find length) length * width = area width = area/length = (x^2-5x+6)cm^2 / ((5x-15)/4) cm = (x^2-5x+6) / ( 5x-15)/(4) cm = (x-2)/20 cm (after some calculations)